# Two Tales of Persona in LLMs: A Survey of Role-Playing and Personalization

### v1

@@ Abstract
Recently, methods investigating how to adapt large language models (LLMs) for specific scenarios have gained great attention. Particularly, the concept of persona, originally adopted in dialogue literature, has re-surged as a promising avenue. However, the growing research on persona is relatively disorganized, lacking a systematic overview. To close the gap, we present a comprehensive survey to categorize the current state of the field. We identify two lines of research, namely (1) LLM Role-Playing, where personas are assigned to LLMs, and (2) LLM Personalization, where LLMs take care of user personas. To the best of our knowledge, we present the first survey tailored for LLM role-playing and LLM personalization under the unified view of persona, including taxonomy, current challenges, and potential directions. To foster future endeavors, we actively maintain a paper collection available to the community. [^1] https://github.com/MiuLab/PersonaLLM-Survey **footnotetext: Equal contribution.
@@ A Survey of Role-Playing and Personalization
Yu-Min Tseng*αβ Yu-Chao Huang*α Teng-Yun Hsiao*α Yu-Ching Hsuα Jia-Yin Fooα Chao-Wei Huangα Yun-Nung Chenα αNational Taiwan University βAcademia Sinica ymtseng@nlg.csie.ntu.edu.tw {R11222015, B10502058}@ntu.edu.tw y.v.chen@ieee.org
Figure 1: An illustrated overview of the survey. In Role-Playing, LLMs act according to assigned personas (i.e., roles) under a defined environment. For example, given role names (red) with descriptions (black), LLMs are tasked to role-play in a social simulation scenario. For Personalization, LLMs consider users’ personas (blue) to generate tailored responses to the same question. The dashed rectangles indicate prompts and the solid rectangles indicate LLMs’ responses.
@@ 1 Introduction
The striking capabilities of large language models (LLMs), exemplified by ChatGPT OpenAI (2022), have significantly advanced the field of natural language processing (NLP; Wei et al., 2023; Madaan et al., 2024; Shinn et al., 2024). Recently, in addition to using LLMs as NLP task solvers or general-purpose chatbots, the question of how to adapt LLMs for specific scenarios has received great attention. To this end, leveraging personas has resurfaced as an ideal lens for adapting LLMs in target scenarios (Chen et al., 2023a, 2024). With personas, LLMs are equipped to generate tailored responses while harnessing the full capabilities. However, the growing literature on persona in the LLM era is relatively disorganized, lacking a unifying overview.
In this paper, we aim to close the gap by offering a comprehensive survey and a systematic categorization of present studies. Specifically, we divide current research into two main directions, namely LLM Role-Playing and LLM Personalization, as illustrated in Figure 1. The primary distinction is that – in role-playing, the persona belongs to the LLM, while in personalization, the persona belongs to the user. Detailed definitions and key design aspects of each are:
  • LLM Role-Playing. LLMs are tasked to play the assigned personas (i.e., roles) and act accordance to environmental feedback. The key aspect is how LLMs adapt to defined environments.
  • LLM Personalization. LLMs are tasked to take care of users’ personas (e.g., background information, or historical behaviors) to meet customized needs. The key aspect is how LLMs adapt to distinct users.
To the best of our knowledge, we present the first survey for LLM role-playing and LLM personalization under the unified view of persona. To foster future endeavors, we actively maintain a paper collection available to the research community. We hope this work could serve as a beginner guide for newcomers and a practical roadmap for current researchers in the field.
Methodology Dataset
@@
for tree= forked edges, draw=edgeColor, thick, font=, fill=tnodeColor, rectangle, rounded corners=4pt, text=textColor, blur shadow=shadow scale=0.95, shadow xshift=.5ex, shadow yshift=-.5ex, shadow opacity=0.25, edge=-, draw=edgeColor, line width=2pt, grow=0, child anchor=west, parent anchor=east, anchor=west, align=center, l sep+=0.3cm, s sep+=0.3cm, , root/.style= fill=Gray!15, font=, rounded corners=6pt, text width=4cm, /tikz/align=center, inner sep=6pt, tnode2_1/.style=fill=Gray!15, font=, text width=9cm, /tikz/align=center, inner sep=6pt, tnode2_2/.style=fill=Gray!15, font=, text width=6cm, /tikz/align=center, inner sep=6pt, tnode2_3_1/.style=fill=Gray!15, font=, text width=5cm, /tikz/align=center, inner sep=6pt, tnode2_3_2/.style=fill=Gray!15, font=, text width=9cm, /tikz/align=center, inner sep=6pt, tnode2_4/.style=fill=Gray!15, font=, text width=8cm, /tikz/align=center, inner sep=4pt, tnode3_1/.style=fill=Gray!15, font=, text width=8cm, /tikz/align=center, inner sep=6pt, tnode3_2/.style=fill=Gray!15, font=, text width=7cm, /tikz/align=center, inner sep=6pt, tnode3_3/.style=fill=Gray!15, font=, text width=8cm, /tikz/align=center, inner sep=6pt, tnode3_4/.style=fill=Gray!15, font=, text width=6cm, /tikz/align=center, inner sep=6pt, tnode4_1/.style=fill=tnodeColor4, font=, text width=11cm, /tikz/align=center, inner sep=6pt, tnode4_2/.style=fill=tnodeColor4, font=, text width=8.2cm, /tikz/align=center, inner sep=6pt, cnode_1/.style= fill=Yellow!15, font=, edge=dotted, text width=16cm, align=left, cnode_2/.style= fill=Yellow!15, font=, edge=dotted, text width=13cm, align=left, cnode2_1/.style=fill=Green!15, font=, edge=dotted, text width=16cm, align=left, cnode2_2/.style=fill=Green!15, font=, edge=dotted, text width=13cm, align=left, cnode3/.style=fill=tnodeColor4, font=, edge=dotted, text width=22cm, align=left, [Taxonomy, root [LLM Personalization Section 3, tnode2_1 [Methods Section 3.2, tnode2_2 [Prompting, tnode2_3_1 [Profile-Augmented, tnode2_4 [ Richardson et al. (2023);
 ONCE (Liu et al., 2023b), fill=Yellow!15, font=, edge=dotted, text width=11cm, align=left] ] [Retrieval-Augmented, tnode2_4 [ WorkSM (Mysore et al., 2023);
 LAMP (Salemi et al., 2024), fill=Green!15, font=, edge=dotted, text width=11cm, align=left] [ Li et al. (2023a);
 PEARL (Mysore et al., 2023), fill=Yellow!15, font=, edge=dotted, text width=11cm, align=left] ] [Vanilla, tnode2_4 [ Dai et al. (2023);
 Christakopoulou et al. (2023);
 BookGPT (Zhiyuli et al., 2023), fill=Yellow!15, font=, edge=dotted, text width=11cm, align=left] ] ] [Retrieval Augmentation, tnode2_3_2 [ PER-CHAT (Wu et al., 2021);
 PersonaKGC (Fu et al., 2022);
 DuLemon (Xu et al., 2022);
 PerLTQA (Du et al., 2024), cnode2_1] [ Wu et al. (2021);  Fu et al. (2022);
 PLATO-LTM (Xu et al., 2022);  Du et al. (2024);
 MemoryBank (Zhong et al., 2024), cnode_1] ] [Fine-Tuning, tnode2_3_2 [ P-RLHF (Li et al., 2024c), cnode_1] ] ] [Tasks Section 3.1, tnode2_2 [Personalized Education, tnode2_3_2 [ HumSum Shehata et al. (2023);
 EduChat (Dan et al., 2023);  Park et al. (2024), cnode_1] ] [Personalized Healthcare, tnode2_3_2 [ MedDialog (Zeng et al., 2020);
 IMCS-21 (Chen et al., 2022b);
 iCliniq (Li et al., 2023f), cnode2_1] [ openCHA (Abbasian et al., 2024a);
 CHA (Abbasian et al., 2024c);
 MaLP (Zhang et al., 2024a);
 HealthLLM (Jin et al., 2024b), cnode_1], ] [Personalized Search, tnode2_3_2 [ Pass et al. (2006), cnode2_1] [ CoPS (Zhou et al., 2024), cnode_1] ] [Recommendation System, tnode2_3_2 [ MovieLens (Harper and Konstan, 2015), cnode2_1] [ PALR (Yang et al., 2023b);  Dai et al. (2023);
 Christakopoulou et al. (2023);
 BookGPT (Zhiyuli et al., 2023);
 ONCE (Liu et al., 2023b), cnode_1] ] [Personalized
Dialogue, fill=Gray!15, font=
, text width=5cm, /tikz/align=center, inner sep=6pt [User Persona
Modeling, fill=Gray!15, font=, text width=6cm, /tikz/align=center, inner sep=4pt [ PersonaChat (Zhang et al., 2018a);
 ConvAI2 PersonaChat (Dinan et al., 2019), cnode2_2] [ $\mathcal{P}^{2}$ (Liu et al., 2020);  Kim et al. (2024b);
 CLV (Tang et al., 2023b), cnode_2] ] [ToD Modeling, fill=Gray!15, font=, text width=6cm, /tikz/align=center, inner sep=4pt [ MutiWoz (Budzianowski et al., 2018);
 SGD (Rastogi et al., 2020), cnode2_2] [ RefGPT (Yang et al., 2023a);
 ProToD (Hu et al., 2023);
 DSP (Li et al., 2024e), cnode_2] ] ] ] ] [LLM Role-Playing Section 2, tnode3_1 [Agentic
Interactions Section 2.3, tnode3_2 [Emergent Behaviors, fill=Gray!15, font=, text width=7cm, /tikz/align=center, inner sep=6pt [ AgentVerse (Chen et al., 2023c);  Fu et al. (2023);
 MedAgents (Tang et al., 2023a);  CAMEL (Li et al., 2024b), fill=Yellow!15, font=
, edge=dotted, text width=18cm, align=left] ] [Schemas, fill=Gray!15, font=, text width=5cm, /tikz/align=center, inner sep=6pt [Multi-Agent, tnode3_4 [ ChatDev (Qian et al., 2023);
 MedAgents (Tang et al., 2023a), fill=Yellow!15, font=
, edge=dotted, text width=12.8cm, align=left] ] [Single-Agent, tnode3_4 [ Voyager (Wang et al., 2023a);
 MindAct (Deng et al., 2024), fill=Yellow!15, font=, edge=dotted, text width=12.8cm, align=left] ] ] ] [Environments Section 2.2, tnode3_2 [General Framework, tnode3_3 [ CAMEL AI Society; CAMEL Code; CAMEL Math;
 CAMEL Science (Li et al., 2024b), fill=Green!15, font=, edge=dotted, text width=17cm, align=left] [ CAMEL (Li et al., 2024b);  SPP (Wang et al., 2023e);
 AgentVerse (Chen et al., 2023c), fill=Yellow!15, font=, edge=dotted, text width=17cm, align=left] ] [LLM as Evaluators, tnode3_3 [ DRPE (Wu et al., 2023b);  ChatEval (Chan et al., 2023), fill=Yellow!15, font=, edge=dotted, text width=17cm, align=left] ] [Medical Application, tnode3_3 [ Wu et al. (2023a);  MedAgents (Tang et al., 2023a), fill=Yellow!15, font=, edge=dotted, text width=17cm, align=left] ] [Game, fill=Gray!15, font=
, text width=5cm, /tikz/align=center, inner sep=6pt [ Generative Agents (Park et al., 2023);
 Humanoid Agent (Wang et al., 2023f);
 Fu et al. (2023);  Voyager (Wang et al., 2023a), fill=Yellow!15, font=
, edge=dotted, text width=20cm, align=left] ] [Web, fill=Gray!15, font=, text width=5cm, /tikz/align=center, inner sep=6pt [ Mind2Web (Deng et al., 2024);  WebArena (Zhou et al., 2023);
 VisualWebArena (Koh et al., 2024);
 VisualWebBench (Liu et al., 2024), fill=Green!15, font=
, edge=dotted, text width=20cm, align=left] [ MindAct (Deng et al., 2024);  CogAgent (Hong et al., 2023b);
 SeeClick (Cheng et al., 2024a), fill=Yellow!15, font=, edge=dotted, text width=20cm, align=left] ] [Software Development, tnode3_3 [ SoftwareDev (Hong et al., 2023a);
 SRDD (Qian et al., 2023), fill=Green!15, font=, edge=dotted, text width=17cm, align=left] [ Dong et al. (2023);  ChatDev (Qian et al., 2023);
 MetaGPT (Hong et al., 2023a), fill=Yellow!15, font=, edge=dotted, text width=17cm, align=left] ] ] ] ]
@@
Overview.
We organize the survey paper as laid out in the taxonomy (Figure 2). Firstly, we introduce LLM role-playing Section 2, including agent-based role-playing of LLMs Section 2.1, various environments LLMs adapt to Section 2.2, and how LLMs interact within the environments as well as emergent behaviors stemmed from their interactions Section 2.3. Subsequently, we elaborate LLM personalization Section 3, outlining the diverse tasks Section 3.1 and methods Section 3.2 for personalizing LLMs. The benchmarks and datasets for various environments and personalized scenarios are introduced in relevant sections accordingly. We also add a separate branch Section 4 for LLM personality evaluation focusing on psychological assessments. Lastly, we highlight current challenges in role-playing and personalized LLMs and suggest potential future directions Section 5.
@@ 2 LLM Role-Playing
Figure 3: An illustration of five LLM role-playing environments: Software Development Section 2.2.1, Web Section 2.2.2, Game Section 2.2.3, Medical Application Section 2.2.4, and LLM as Evaluators Section 2.2.5. For each environment, we provide a simple scenario with a task description (red-bordered) and relevant personas (i.e., roles; blue-bordered). The dashed rectangle represents an example LLM role-playing prompt template. In addition to the above environments, several research also proposes general frameworks applicable to different environments Section 2.2.6.
In order to perform diverse tasks effectively, a potential solution is to train specialized LLMs tailored to each task. However, this process is computationally prohibitive and time-consuming. To address this, recent studies have demonstrated that leveraging role-playing increases LLMs capabilities, including reasoning (Wu et al., 2023a; Kong et al., 2023), planning Huang et al. (2024); Li et al. (2024b), and problem-solving (Wang et al., 2023e; Qian et al., 2023). Specifically, role-playing triggers the corresponding inherent personalities of LLMs to generate responses aligned with given roles and environments. Equipped with assigned personas (i.e., roles), LLMs adapt adequately to defined environments, thereby enhancing their performance.
In this section, we first clarify the relationship between LLM role-playing and LLM-based agent  Section 2.1. Next, we introduce diverse environments that LLMs could adapt to Section 2.2. Lastly, we identify agentic interactions between LLMs within defined environments and emergent behaviors stemming from their interactions Section 2.3.
2.1 LLM-based Agent
Recently, several research has explored the planning, reflection, and tool-using abilities in LLMs, potentially paving the way for artificial general intelligence (AGI; Bubeck et al., 2023). Thus, LLM-based agent has surged as an fast-growing research field (Wang et al., 2024b; Xi et al., 2023). Moreover, the Multi-Agent framework, where multiple LLM-based agents cooperate and communicate with each other in order to solve complex tasks, has also appeared as a prevalent research direction (Guo et al., 2024).
To answer the key question discussed in Section 1: how LLMs adapt to defined environments, the predominant approach is LLM-based agents with role-playing, where LLMs act according to the assigned personas to interact with the environments. For instance, one of the first works of LLM-based agents proposes generative agents (Park et al., 2023), which adopts role-playing to mimic human behaviors. By giving a list of sentences with name, age, and personality trait that specifies the role, generative agents can perform accordingly and engage in a social simulation environment.
@@ 2.2 Environments
One key aspect of the agentic frameworks is the environments, within which the LLMs could operate and interact given certain rules and receive feedback. We classify the environments for LLM role-playing into five categories: software development, web, game, medical application, and LLM as evaluators, detailed in Section 2.2.1 to Section 2.2.5. Furthermore, we explore a range of frameworks that generalize across different environments in Section 2.2.6.  Figure 3 demonstrates an illustration by providing a simple example task of each environment.
@@ 2.2.1 Software Development
The tasks of software development are to design programs, software, or coding projects. Example tasks include “Create a snake game.”, “Create a Python program to develop an interactive weather dashboard.” (Hong et al., 2023a), and HumanEval (Chen et al., 2021). Due to the complexity of these tasks, often too intricate to be completed correctly on the first attempt, existing research predominantly leverages the Waterfall model Petersen et al. (2009); Bassil (2012) or Standardized Operating Procedures (SOPs) Belbin and Brown (2022); DeMarco and Lister (2013) to break down the tasks into manageable sub-tasks.
@@
Similar to real-world software development projects, a collaborative multi-agent framework with LLM role-playing operating as a company serves as an effective approach for software development environment Qian et al. (2023); Hong et al. (2023a); Dong et al. (2023). Concretely, the roles might include Chief Technology Officer (CTO), Chief Product Officer (CPO), Chief Executive Officer (CEO), Product Manager, Engineer, Reviewer, and Tester. By assigning specific roles, LLMs are capable of generating content in a step-by-step and accurate manner.

### v2

@@ Abstract
The concept of persona, originally adopted in dialogue literature, has re-surged as a promising framework for tailoring large language models (LLMs) to specific context (e.g., personalized search, LLM-as-a-judge). However, the growing research on leveraging persona in LLMs is relatively disorganized and lacks a systematic taxonomy. To close the gap, we present a comprehensive survey to categorize the current state of the field. We identify two lines of research, namely (1) LLM Role-Playing, where personas are assigned to LLMs, and (2) LLM Personalization, where LLMs take care of user personas. Additionally, we introduce existing methods for LLM personality evaluation. To the best of our knowledge, we present the first survey for role-playing and personalization in LLMs under the unified view of persona. We continuously maintain a paper collection to foster future endeavors. **footnotetext: Equal contribution.
@@ A Survey of Role-Playing and Personalization
[…]
Figure 1: In Role-Playing, LLMs act according to assigned personas (i.e., roles) under a defined environment. For example, given role names with descriptions, LLMs role-play in a social simulation game. For Personalization, LLMs consider user personas to generate tailored responses to the same question. Dashed rectangles are prompts and solid rectangles are LLMs’ responses.
@@ 1 Introduction
The striking capabilities of large language models (LLMs), exemplified by ChatGPT OpenAI (2022), have significantly advanced the field of natural language processing (NLP; Wei et al., 2023; Madaan et al., 2024; Shinn et al., 2024). Recently, in addition to using LLMs as NLP task solvers or general-purpose chatbots, the question of how to adapt LLMs for specific context has received great attention. To this end, leveraging personas has resurfaced as an ideal lens for adapting LLMs in target scenarios (Chen et al., 2023a, 2024). By incorporating personas, LLMs can generate more contextually appropriate responses, maximizing their utility and effectiveness for specific applications. However, the growing literature on persona in the LLM era is relatively disorganized, lacking a unifying overview.
@@
for tree= forked edges, draw=edgeColor, thick, font=, fill=tnodeColor, rectangle, rounded corners=4pt, text=textColor, blur shadow=shadow scale=0.95, shadow xshift=.5ex, shadow yshift=-.5ex, shadow opacity=0.25, edge=-, draw=edgeColor, line width=2pt, grow=0, child anchor=west, parent anchor=east, anchor=west, align=center, l sep+=0.3cm, s sep+=0.1cm, , root/.style= fill=Gray!15, font=, rounded corners=6pt, text width=4cm, /tikz/align=center, inner sep=6pt, tnode2_1/.style=fill=Gray!15, font=, text width=12cm, /tikz/align=center, inner sep=6pt, tnode2_2/.style=fill=Gray!15, font=, text width=11cm, /tikz/align=center, inner sep=6pt, tnode2_3/.style=fill=Gray!15, font=, text width=12cm, /tikz/align=center, inner sep=6pt, tnode3_1/.style=fill=Gray!15, font=, text width=12cm, /tikz/align=center, inner sep=6pt, tnode3_2/.style=fill=Gray!15, font=, text width=11cm, /tikz/align=center, inner sep=6pt, tnode3_3/.style=fill=Gray!15, font=, text width=12cm, /tikz/align=center, inner sep=6pt, [Taxonomy, root [LLM Personality Evaluation Section 4, tnode2_1 [BigFive; MBTI; etc, tnode2_2] ] [LLM Personalization Section 3, tnode2_1 [Dialogue Section 3.5, tnode2_2 [User Persona Modeling, tnode2_3] [Task-Oriented Modeling, tnode2_3] ] [Healthcare Section 3.4, tnode2_2] [Education Section 3.3, tnode2_2] [Search Section 3.2, tnode2_2] [Recommendation Section 3.1, tnode2_2] ] [LLM Role-Playing Section 2, tnode3_1 [Emergent Behaviors Section 2.3, tnode3_2 [Destructive Behavior, tnode3_3] [Conformity Behavior, tnode3_3] [Voluntary Behavior, tnode3_3] ] [Role-Playing Schema Section 2.2, tnode3_2 [Multi-Agent, tnode3_3] [Single-Agent, tnode3_3] ] [Environments Section 2.1, tnode3_2 [LLM-as-Evaluator Section 2.1.4, tnode3_3] [Medical Application Section 2.1.3, tnode3_3] [Game Section 2.1.2, tnode3_3] [Software Development Section 2.1.1, tnode3_3] ] ] ]
@@
In this paper, we aim to close the gap by offering a comprehensive survey and a systematic categorization of existing studies. Specifically, we divide current research into two main streams, namely LLM Role-Playing and LLM Personalization, as illustrated in Figure 1. The primary distinction is that in role-playing, the persona belongs to the LLM, while in personalization, the persona belongs to the user. The definitions are detailed below.
  • LLM Role-Playing: LLMs are tasked to play the assigned personas (i.e., roles) and act based on environmental feedback, adapting to the environment.
  • LLM Personalization: LLMs are tasked to take care of user personas (e.g., background information or historical behaviors) to meet individualized needs, adapting to distinct users.
To the best of our knowledge, we present the first survey for LLM role-playing and LLM personalization under the unified view of persona. To foster future endeavors, we actively maintain a paper collection available to the research community. We aim for this work to serve as both a valuable introduction for newcomers and a comprehensive resource for current researchers in the field.
Our taxonomy is illustrated in Figure 2. We first introduce LLM role-playing Section 2, followed by LLM personalization Section 3. Next, we provide an overview of evaluation methods Section 4 assessing whether the personality of LLMs (e.g., personality traits or psychological behaviors) accurately aligns with expected personas after the adaptation (i.e., for role-playing LLMs that act according to assigned personas and personalized LLMs that fit user personas). Lastly, we highlight current challenges and future directions Section 5. A comprehensive list of benchmarks and datasets is provided in the Appendix.
@@ 2 LLM Role-Playing
Figure 3: An illustration of five LLM role-playing environments: Software Development Section 2.1.1, Game Section 2.1.2, Medical Application Section 2.1.3, and LLM as Evaluators Section 2.1.4. For each environment, we provide a simple scenario with a task description (red-bordered) and relevant personas (i.e., roles; blue-bordered). The dashed rectangle represents an example LLM role-playing prompt template. In addition to the above environments, past research also proposes general frameworks applicable to different environments Section 5.1.
LLM-based language agents have demonstrated impressive abilities, such as planning, reflection, and tool-use recently Yao et al. (2022b); Shinn et al. (2024); Yao et al. (2024). The predominant approach of LLM role-playing is by coupling personas with language agents, specifically, by incorporating personas directly inside the prompt of language agents. Such a training-free paradigm is particularly desirable due to its simplicity and effectiveness.
Language agents with role-playing elicit the corresponding parametric knowledge in LLMs to generate responses aligned with assigned personas (i.e., role), enabling them to adapt to various interactive environments. LLM role-playing also extends to multi-agent settings, where multiple language agents are equipped with diverse personas, cooperating and communicating with each other to solve complex tasks (Guo et al., 2024). For instance, in one of the first works of role-played LLMs, Park et al. (2023) propose generative agents, which engage in a social simulation environment by mimicking human behaviors according to names, ages, and personality traits specified in the prompts.
Following we introduce different environments and associated roles in which LLMs adapt to Section 2.1, interactions between LLMs within the environment Section 2.2, and emergent behaviors stemming from their interactions Section 2.3. Figure 3 provides an illustrative overview.
@@ 2.1.1 Software Development
For software development, the goal typically involves designing programs or coding projects. For instance, “Create a snake game.” or “Create a Python program to develop an interactive weather dashboard.” (Hong et al., 2023a). Due to the complexity of these tasks, often too intricate to be completed correctly on the first attempt, existing research leverages approaches like the Waterfall model Petersen et al. (2009); Bassil (2012) or Standardized Operating Procedures (SOPs) Belbin and Brown (2022); DeMarco and Lister (2013) to break down the tasks into manageable sub-tasks.
@@
Similar to real-world settings, LLMs role-play to operate as a company in a collaborative, multi-agent software development environment Qian et al. (2023); Hong et al. (2023a); Dong et al. (2023). Different roles include Chief Technology Officer (CTO), Chief Product Officer (CPO), Chief Executive Officer (CEO), Product Managers, Engineers, Reviewers, and Testers. By assigning specific roles, LLMs are capable of carrying out tasks in a step-by-step and accurate manner.